OpenStudy (anonymous):

If 7 is added to twice a number and this sum is multiplied by 4, the result is the same as if the number is multiplied by 4 and 8 is added to the product. What is the number?

OpenStudy (tanner23456):

Set up the equation (7 + 2x) 4 = (x*4) +8 solve for x

OpenStudy (anonymous):

Alright the answer I got is 1/4 is this right

OpenStudy (tanner23456):

Multiply the 4 in 28 + 8x = 4x +8 Subtract the 4x 28 + 4x = 8 Subtract the 28 4x = -20 divide the 4 x = -5

OpenStudy (anonymous):

Alright I see the step I did wrong thanks for your help though

OpenStudy (tanner23456):

No worries